# UNet with Axial Transformer : A Neural Weather Model for Precipitation Nowcasting

### v1

@@ 4 Experimental Results
In this section, we first introduce related datasets and implementation details in our experiments. Then, we report results and comparison of all the experiments conducted on each of the model used.
@@ 4.1 Datasets
The dataset used in this project is a part of dataset used in paper ConvLSTM for Precipitaion Nowcasting [14]. The dataset for the original research work consisted of HKO-7 dataset which contains radar echo data from 2009 to 2015 near Hong Kong. Along with the radar images, the dataset also compressed of proportions of rainfall events with different rain-rate threshold.
The HKO-7 dataset used in the baseline research contains radar CAPPI reflectivity images, which have resolution of 480×480 pixels, are taken from an altitude of 2km and cover a 512km × 512km area centered in Hong Kong. The data are recorded every 6 minutes and hence there are 240 frames per day. The pixel values are clipped between 0 and 255. There are some noisy radar images in the dataset generated by factors like ground clutter, sea clutter, anomalous propagation and electromagnetic interference.
The complete dataset with all the csv and pkl files was not available publicly. But a part of it, i.e., the radar images that can be conveniently used in our problem are available publicly [5]. We have continuous radar images from 514 folders. Each folder contains variable number of radar images sampled, but all of them are continuous. As the radar echo maps arrive in a stream, nowcasting algorithms can apply online learning to adapt to the newly emerging spatiotemporal patterns.
This was the most convenient dataset we could use, because, the datasets used in previous research under this topic were too large to be handled on the GPU provided (datasets of size in TBs). Other dataset were either paid or not publicly available. Even the dataset that we have used, is a part of a original dataset. But it was enough to carry out several experimentations on various models used.
@@ 4.2 Implementation Details
After taking the dataset as our input, first we had to ensure that the number of radar images in each folder’s sequence remain same. This is due to the fact that we will be sending the dataset in batches for training and testing and we need to have uniform number of images in each batch. Hence, we clipped the dataset to 240 images per folder (if the number of radar images in a folder is greater than 240, if it is less than 240 we reject that folder), i.e. as soon as the number of images in that particular folder go past 240, we stop taking that folder’s images as input. We found out that dataset now had 482 sequences of length 240. While we input each image, we also convert the image into grayscale, as we need to map the cloud motion in the radar images and having RGB channels in the images will only cost computation time and power without much significant difference in the results.
At first, we also tried to send these data directly for training in batches. But again, due to computational power limitations, we could not send a batch of size 240 for training. The sequence of images again had to be broken to the length of 20, i.e. instead of 482 sequences of length 240, now we had 5784 sequences of length 20. Now sequence of this length can be comfortably loaded for training and testing.
Now that we had sequence of data, we also needed to filter out the images that are either too dark or too light. There were some faulty images in our dataset as previously talked about because of faulty radar images. So we first stored the summation of all the pixel values in each image of our dataset in a list. This is important to normalise our dataset, as then we found the 25th and 75th percentile of the list values. These gives us the range we will use to filter images whose sum of pixel values are above or below the specified range.
Now going through each sequence, we again calculated the sum of pixel values of each image in 5784 sequences, if the sum is outside the range, we count that as a bad-frame. If the number of bad-frames in a sequence exceed 10, we discard that sequence as a whole, as we cant discard few frames from a sequence; discontinuity in frames would make them useless. Now once the data is cleaned, we are now left with 2901 sequences of length 20, each being an image of size 128 x 128. The sequences then were loaded in an .npy file to used later. As observed the usable images left at last very almost half of the dataset we created at start and this was one of the reasons why we couldnt use higher batch size while training our models.
@@
Configuration ConvLSTM cGANs UNet
Encoder-Decoder
with Axial Transformer
Table 1: While testing each of the model, we first input sequence (length similar to that used for training of each model) of images ($t_{0},t_{1},...,t_{m}$) into the model and get 1 image as output ($p_{1}$) from it. Now this predicted image $p_{1}$ is used along with input sequence ($t_{1},..,t_{m}$) to predict the next radar image and so on. The result of each model are given at the top and as we can see the novel model has the highest quality of prediction.
Once the .npy file is loaded, we randomly shuffled these sequences without altering their internal continuity in frames, for uniform distribution of types of images for training, validation and testing. Out of these, 2000 sequences were used in training dataset, 500 for validation and 400 for testing purpose. Whenever the numpy array representing the images in the dataset is loaded, the batch is is normalised by dividing by 255, to bring each pixel value between 0 and 1. Shuffle is kept true to make the distribution among the types of images uniform.
The sequence length given to train Conv-LSTM was 15, so that it can count next 5 frames as target and make prediction on those. For cGANs, the input sequence length was 4, as it was an image-to-image translation technique and produced one image per input image, hence predicting 4 frames from 4 inputs. For UNet and UNet with axial attention, we had to give 16 frames as input to the model as we needed to ensure downsampling and later upsampling produce images (given in number of channels dimensions) of same number at each corresponding layer. The batch size for ConvLSTM = 2, cGANs = 4, UNet = 4, UNet with axial transformer = 1. Every model was given a batch size with limitation due to risk of GPU running out of memory or model overfitting too soon.
For a given input sequence, 1 frame was predicted as output from every model. Now the same frame was used as input while predicting the next frame and so on. Hence, by this method we were able to predict 4-5 frames in future using the original sequence of input image.
Number of epochs used to train the models is variable due to effect of overfitting for larger and deeper models. ConvLSTM: 5, cGANs: 10, UNet: 5, UNet with Axial Transformer: 15. Optimizer used is Adam, with a learning rate of: 1e-3 for ConvLSTM, 1e-4 for cGANs, 1e-3 for UNet, 1e-4 for UNet with Axial Transformer. The loss criterion for each of the model used is MSELoss.
Given a noise-free flattened monochrome image I and its flattened noisy approximation Image K, MSELoss between two tensors is defined as:
$MSELoss\ =\ \frac{1}{2}\sum_{i=0}^{n-1}(\ I_{i}\ -\ K_{i}\ )$ (7)
The metrices we used to compare the results of different models were: PSNR and SSIM. PSNR: Peak signal-to-noise ratio is the ratio between the maximum possible power of a signal and the power of corrupting noise that affects the fidelity of its representation.
@@
MSE is defined as mean squared error between a noise-free flattened monochrome image and its flattened noisy approximation Image K
@@
$SSIM(\ x,y)\ =\ \frac{(2u_{x}u_{y}+c_{1})(2\sigma_{xy}+c_{2})}{\left(u_{x}^{2}+u_{y}^{2}+c_{1}\right)\left(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2}\right)}$ (9)
@@ 4.3 Results and Comparisons
While training each of the model we calculated the loss for training and cross validation data, the results of which are given below:
[Uncaptioned image][Uncaptioned image]
Figure 2: cGANs training
[Uncaptioned image][Uncaptioned image]
Figure 3: UNet training
[Uncaptioned image][Uncaptioned image]
Figure 4: UNet with Axial Transformer training
We see some unusual spikes in the graph which might be due to faulty radar images and abrupt shuffling of data.
The models used in the experiments showed a gradual increase in both the metrices and hence show that our novel architecture performs better than the baseline model used.
Configuration PSNR SSIM
@@
Encoder-Decoder
Axial Transformer
47.6678 0.9943
Table 2: The table shows comparison of results from each model on the test dataset
While the baseline ConvLSTM showed satisfactory results on the test dataset there was an improvement when cGANs were used. Using the pixel to pixel translation in radar images using generator and discriminator, the model is robust to random noise introduced.
The generator used in cGANs was UNet which was able to give much better results while it was used alone. This might be possible due to the fact that dataset used was comparatively small to be used by larger and deeper models. Hence, a simple model can perform better when there is less complexity involved in learning. Furthermore, the main task of GANs model is to generate new images with same statitics, hence with pix2pix translation we could only use a single image to generate another image which corresponds to a sequence of image to generate other half of sequence and hence almost half the dataset available for training as compared to rest of the models.
The skip connections in UNet not only helps to resolve vanishing gradient problem, but also helps in feature reusability. With a very simple architecture it is able to embed enough information of sequence to predict the next frame in the sequence. A further enhancement was use of axial transformers on top of UNet. This ensured that model is able to attend the pixels in same rows and columns making it more suitable for trajectory tracking problems like these.
While testing a model, once we have an output, we actual have its pixel value scaled between 0 and 1, due to initial normalisation applied on input data. Hence, we first see if the value of pixel is above some threshold (typically 0.2), if it is, we multiply it with 255, hence directly converting it to a white pixel. This is the reason why we have B&W image in the output.
@@ 5 Discussion
In this paper, we propose a novel framework for precipitation nowcasting by predicting 4-5 frames in future using a sequence of frames from past, each 6 minutes apart. Our models exploits spatio-temporal features of the radar images and create realistic predictions. Moving beyond traditional weather forecasting we exhibit state-of-art solution for the HKO-7 dataset used.
We experimented on several models to find the best fit for the problem and found that UNet paired with Axial Transformer is able to predict the next radar images more accurately than any of the other models used, due to the skip connections and ability to attend pixels in same row and column. We started with a baseline model and added components systematically and saw the results amplifying.
With enhanced dataset like that including actual rainfall values with images can help in building end-to-end model to predict % rainfall for a given area. The methodology used can also find its use in the areas of traffic management, bio-medical, and autonomous vehicles, due to being able to counter the fundamental problem of next frame prediction.
@@ 6 Future Work
As discussed earlier, the pixel values of radar images given as input to our model was normalised between 0 and 1. Currently the goal of the project was to predict the radar images that match the trajectory of the current motion of the clouds in the given sequence of images, hence in the output we essentially classify each pixel as cloud or not cloud. In future, we could keep the pixel values in images as it is and hence calculate the amount of precipitation under each pixel using the pixel value. Higher the pixel value, higher the intensity of precipitation.
We also converted the radar images into GRAYSCALE during the preprocessing this was important to reduce the computational cost but with little to no effect in results due to original images themselves being B&W colour. If we are able to work on dataset that consists of colourful satellite images with RGB channels, first we could do image segmentation in the radar images identifying the chunk of land and cloud as seprate and predicting the precipitation intensity corresponding to each of the location in the radar images. This location can be town, city or even as huge as states as they are mapped from actual satellite images and not radar images.

### v2

@@ 4 Experimental Results
In this section, we first introduce related datasets and implementation details in our experiments. Then, we report results and a comparison of all the experiments conducted on each of the models used.
@@ 4.1 Datasets
The dataset used in this project is a part of dataset used in paper ConvLSTM for Precipitation Nowcasting [18]. The dataset for the original research work consisted of HKO-7 dataset, which contains radar echo data from 2009 to 2015 near Hong Kong. Along with the radar images, the dataset also contains proportions of rainfall events with different rain-rate thresholds.
The HKO-7 dataset used in the baseline research contains radar CAPPI reflectivity images, which have a resolution of 480×480 pixels, are taken from an altitude of 2km, and cover a 512km × 512km area centered in Hong Kong. The data are recorded every 6 minutes, and hence there are 240 frames per day. The pixel values are clipped between 0 and 255. There are some noisy radar images in the dataset generated by factors like ground clutter, sea clutter, anomalous propagation, and electromagnetic interference.
The complete dataset with all the CSV and pkl files was not available publicly. But a part of it, i.e., the radar images that can be conveniently used in our problem, are available publicly [8]. We have continuous radar images from 514 folders. Each folder contains a variable number of radar images sampled, but all of them are continuous. As the radar echo maps arrive in a stream, nowcasting algorithms can apply online learning to adapt to the newly emerging spatiotemporal patterns.
This was the most convenient dataset we could use, because the datasets used in previous research under this topic were too large to be handled on the GPU provided (datasets of size in TBs). Other datasets were either paid or not publicly available. Even the dataset that we have used, is a part of an original dataset. But it was enough to carry out several experiments on various models used.
@@ 4.2 Implementation Details
After taking the dataset as our input, first, we had to ensure that the number of radar images in each folder’s sequence remains the same. This is due to the fact that we will be sending the dataset in batches for training and testing, and we need to have a uniform number of images in each batch. Hence, we clipped the dataset to 240 images per folder (if the number of radar images in a folder is greater than 240, if it is less than 240, we reject that folder), i.e., as soon as the number of images in that particular folder goes past 240, we stop taking that folder’s images as input. We found out that the dataset now had 482 sequences of length 240. While we input each image, we also convert the image into grayscale, as we need to map the cloud motion in the radar images, and having RGB channels in the images will only cost computation time and power without much significant difference in the results.
At first, we tried to send this data directly for training in batches. But, due to computational power limitations, we could not send a batch of size 240 for training. The sequence of images had to be broken to the length of 20, i.e., instead of 482 sequences of length 240, now we had 5784 sequences of length 20. Now sequence of this length can be comfortably loaded for training and testing.
Now that we had a sequence of data, we also needed to filter out the images that were either too dark or too light. There were some faulty images in our dataset, as previously talked about, because of faulty radar images. So we first stored the summation of all the pixel values in each image of our dataset in a list. This is important to normalise our dataset, as then we found the 25th and 75th percentile of the list values. These give us the range we will use to filter images whose sum of pixel values is above or below the specified range.
Now, going through each sequence, we again calculated the sum of pixel values of each image in 5784 sequences. If the sum is outside the range, we count that as a bad frame. If the number of bad frames in a sequence exceeds 10, we discard that sequence as a whole, as we can’t discard a few frames from a sequence; discontinuity in frames would make them useless. Now, once the data is cleaned, we are left with 2901 sequences of length 20, each being an image of size 128 x 128. The sequences were then loaded in an .npy file to be used later. As observed, the usable images left at last were almost half of the dataset we created at the start, and this was one of the reasons why we couldn’t use a higher batch size while training our models.
@@
Configuration ConvLSTM cGANs UNet Encoder-Decoder with Axial Transformer
Table 1: While testing each model, we first input a sequence of images ($t_{0},t_{1},...,t_{m}$) into the model and obtain one image ($p_{1}$) as output. This predicted image $p_{1}$ is then used with the next sequence ($t_{1},...,t_{m}$) to predict the next radar image, and so on. The top row shows the target images, and the second row shows the outputs. The proposed model achieves the highest quality of prediction.
Once the .npy file is loaded, we randomly shuffle these sequences without altering their internal continuity in frames, for uniform distribution of types of images for training, validation, and testing. Out of these, 2000 sequences were used in training dataset, 500 for validation, and 400 for testing purpose. Whenever the numpy array representing the images in the dataset is loaded, the batch is normalised by dividing by 255, to bring each pixel value between 0 and 1. Shuffle is kept true to make the distribution among the types of images uniform.
The sequence length given to train Conv-LSTM was 15, so that it can count next 5 frames as target and make a prediction on those. For cGANs, the input sequence length was 4, as it was an image-to-image translation technique and produced one image per input image, hence predicting 4 frames from 4 inputs. For UNet and UNet with axial attention, we had to give 16 frames as input to the model, as we needed to ensure downsampling and later upsampling produce images (given in number of channels dimensions) of the same number at each corresponding layer. The batch size for ConvLSTM = 2, cGANs = 4, UNet = 4, Axial-UNet = 1. Every model was given a batch size with limitations due to the risk of the GPU running out of memory or model overfitting too soon.
For a given input sequence, 1 frame was predicted as output from every model. Now the same frame was used as input while predicting the next frame and so on. Hence, by this method, we were able to predict 4–5 frames in the future using the original sequence of the input image.
Number of epochs used to train the models is variable due to the effect of overfitting for larger and deeper models. ConvLSTM: 5, cGANs: 10, UNet: 5, Axial-UNet: 15. Optimizer used is Adam, with a learning rate of: 1e-3 for ConvLSTM, 1e-4 for cGANs, 1e-3 for UNet, 1e-4 for Axial-UNet. The loss criterion for each of the models used is MSELoss.
Given a noise-free flattened monochrome image I and its flattened noisy approximation Image K, the MSELoss between two tensors is defined as:
[…]
The metrics we used to compare the results of different models were PSNR and SSIM. PSNR: Peak signal-to-noise ratio is the ratio between the maximum possible power of a signal and the power of corrupting noise that affects the fidelity of its representation.
@@
MSE is defined as the mean squared error between a noise-free, flattened monochrome image and its flattened noisy approximation, Image K
@@
[…]
@@ 4.3 Results and Comparisons
We limit evaluation to PSNR/SSIM. Event/object-based skill and cell-tracking metrics are standard in meteorology [13]; comparison to recent diffusion/transformer models is left for future work. While training each of the model, we calculated the loss for training and cross-validation data, the results of which are given below:
(a) cGANs training loss
(b) cGANs validation loss
(c) UNet training loss
(d) UNet validation loss
(e) UNet+Transformer training loss
(f) UNet+Transformer validation loss
Figure 2: Training and validation loss curves for (a,b) cGANs, (c,d) UNet, and (e,f) Axial-UNet.
We see some unusual spikes in the graph, which might be due to faulty radar images and abrupt shuffling of data.
The models used in the experiments showed a gradual increase in both metrics, and hence show that our proposed architecture performs better than the baseline model used.
Table 2: Comparison on the HKO-7 subset. Higher is better.
[…]
@@
Encoder-Decoder with Axial Transformer 47.6678 0.9943
While the ConvLSTM baseline performed reasonably on our split, a pix2pix-style cGAN improved over it. Treating the task as image-to-image translation (one output frame per input frame) with an adversarial loss helps suppress some noise and sharpen texture compared to pure regression.
The cGAN’s generator is a UNet. Interestingly, the same UNet used as a standalone deterministic model performed even better, likely because our dataset is relatively small and the supervised loss dominates; simpler models can generalize well in this regime. In addition, pix2pix yields fewer distinct training pairs than sequence-to-sequence setups, since each training example maps a single input image to a single target image.
Skip connections in UNet help alleviate vanishing gradients and promote feature reuse, enabling accurate next-frame predictions with a compact backbone. Adding axial-attention blocks on top of UNet further improves quality by attending along rows and columns to capture large-scale advection while keeping memory modest.
For evaluation, the model outputs continuous predictions in $[0,1]$ (due to input normalization), and PSNR/SSIM are computed on these continuous fields. Thresholded black-and-white visualizations (e.g., applying a 0.2 cutoff and rescaling to 0–255) are used only for display, not for metric computation.
@@ 5 Discussion
We studied short–lead-time radar nowcasting with a lightweight encoder–decoder augmented by axial attention. Given $M$ past frames (here M=16M{=}16) sampled every 6 minutes, the model predicts the next frame and extends to 4–5 future frames via autoregression. On our HKO-7 subset, the axial-attention variant improves pixel-fidelity metrics (PSNR/SSIM) over ConvLSTM, pix2pix-style cGANs, and a plain UNet. Axial attention preserves coherent structures and sharper boundaries aligned with large-scale advection while keeping memory overhead modest by attending along rows and columns instead of full 2D self-attention.
This study has practical constraints: evaluation is limited to PSNR/SSIM (image fidelity) rather than meteorological skill; a fuller assessment would include event/object-based metrics and lead-time breakdowns. The data are a cleaned, grayscale subset of HKO-7 at $128{\times}128$, which may bias case difficulty; we do not convert reflectivity to rain rate nor assess calibration/uncertainty (we take the mean of the predictive distribution). Baselines follow customary input lengths—ConvLSTM (15), cGAN (4), UNet/ours (16)—and their usual training settings, so the comparison is pragmatic rather than strictly matched-budget.
@@ 6 Future Work
We plan to extend the study along four axes:
  • Continuous targets and rain rate. Train on reflectivity in dBZ and predict continuous fields; convert to rain rate via an appropriate ZZRR relation, and evaluate thresholded rain events rather than binarized visualizations.
  • Meteorology-grade evaluation. Beyond PSNR/SSIM, report CSI, POD, FAR, FSS, and cell-tracking growth/decay across lead times (e.g., 6–30 min).
  • Probabilistic outputs. Retain the distributional head, assess calibration (e.g., CRPS, reliability diagrams), and sample ensembles for uncertainty-aware nowcasts.
  • Data & baselines. Use storm- or date-based splits to avoid leakage; quantify cleaning/clutter effects; add persistence/optical-flow controls and stronger deep baselines under the same split; explore higher resolution and multi-channel inputs (multi-elevation radar, satellite bands).